# Sparse Auto-Encoder Interprets Linguistic Features in Large Language Models

### v1

@@ Appendix A Dataset Construction
A.1 Linguistic Structure Definition
@@
    Phonetics examines the physical production and acoustic properties of speech sounds.
@@
    Phonology investigates the abstract sound systems and patterns in a language.
@@
    Morphology focuses on the internal structure of words.
@@
    Syntax deals with sentence structure and the rules governing the arrangement of words into phrases and clauses.
  5. 5.
    Semantics explores the meaning of words and sentences at a literal or denotational level.
  6. 6.
    Pragmatics considers how context influences meaning, covering phenomena like implicature, presupposition, and speech acts.
A.2 Construction Process
When constructing the dataset, we first manually create 5–10 Chinese and English sentences that strongly exhibit the target linguistic feature, along with 5–10 counterfactual sentences. Next, we use GPT-o1 to generate the remaining portion of the dataset, which includes 80 feature-consistent Chinese and English sentences and 80 counterfactual sentences. Minimal pairs are constructed by manually modifying each feature-consistent sentence. After dataset construction, all sentences are manually reviewed and corrected.
A.3 Construction Examples
We present three examples from morphology, syntax, and semantics.
Plural Noun – Minimal Pairs: Directly change the plural form to singular, disregarding grammar.
The books on the shelf are all about history.
The book on the shelf are all about history.
She bought several flowers to decorate the house.
She bought several flower to decorate the house.
Plural Noun – Counterfactual Sentence: Maintain grammatical correctness while removing the plural form.
The books on the shelf are all about history.
The book on the shelf is all about history.
She bought several flowers to decorate the house.
She bought a flower to decorate the house.
Genitive – Minimal Pair: Remove “of” while keeping the rest unchanged.
The roof of the house was damaged in the storm.
The roof the house was damaged in the storm.
The color of the sky changed at sunset.
The color the sky changed at sunset.
Genitive – Counterfactual Sentence: Retain “of” but use a non-genitive context.
The roof of the house was damaged in the storm.
She is afraid of spiders.
The color of the sky changed at sunset.
He ran out of time.
Contrast – Minimal Pair: Remove the contrast marker, disregarding grammar.
I wanted to go out, but it was raining.
I wanted to go out, it was raining.
She was very tired, yet she kept working.
She was very tired, she kept working.
Contrast – Counterfactual Sentence: Change the meaning by altering the logical relation in the second clause to a continuation.
I wanted to go out, but it was raining.
I wanted to go out, then I grabbed an umbrella.
She was very tired, yet she kept working.
She was very tired, then she took a short nap.
A.4 Special Cases for Minimal Pairs
In cases involving transitive verbs, intransitive verbs, linking verbs, and preposed verbs in inversion, direct deletion results in sentences that lose their predicate meaning and cannot convey a complete semantic unit. In such cases, examining activations is meaningless, and minimal contrast pair datasets are not constructed for these features. In the polite speech dataset, the minimal contrast pairs obtained by removing the politeness marker are identical to the counterfactual sentences converted to non-polite sentences; hence, the two datasets are the same.
Appendix B Feature Extraction
When extracting features using datasets, we employed the following techniques to improve efficiency and accuracy:
B.1 Extraction Modes
We tested the following three analysis modes during feature extraction on the artificially constructed dataset:
Token-Frequency: The total activation frequency of the basis vector across all tokens in the artificial dataset, sorted in descending order.
Sentence-Frequency: The percentage of sentences in the artificial dataset where the basis vector is activated, sorted in descending order.
Sum-Activation: The total activation value of the basis vector across all tokens in the artificial dataset, sorted in descending order.
We build a background dataset with diverse syntactic and semantic patterns. We extract basis vectors that frequently activate across all patterns; their meanings are irrelevant and treated as background noise. They can be removed in our experiments.
We test the analysis modes on the three linguistic features—simile, metaphor, and politeness—identified at layer 26. And we extract 42 background noise basis vectors at layer 26.
To evaluate the efficiency of the three analysis modes—token-frequency, sentence-frequency, and sum-activation—we examine the ranking of target basis vectors in the extracted vectors under each analysis mode. We compare the results both with and without the removal of background noise. The results are presented in Table 4. Based on experimental results, performing frequency analysis at the sentence level and removing background noise is most efficient, as it minimizes the impact of irrelevant or outlier activations on the analysis outcomes.
Modes Simile Metaphor Politeness
T-Freq 99 92 37
T-Freq-NN 66 59 26
S-Freq 29 87 12
S-Freq-NN 13 50 6
S-Act 1135 1257 115
S-Act-NN 1121 1242 111
Table 4: Ranking results of different analysis modes: Token Frequency, Token Frequency without noise, Sentence Frequency, Sentence Frequency without noise, Sum Activation, Sum Activation without noise.
B.2 “20-mix-2” Dataset for Extraction
We first built a dataset composed entirely of sentences highly related to the target linguistic feature. Although directly analyzing sentence-level activation frequencies and sorting the results is feasible, we adopted a more efficient method by inserting 1–2 counterfactual sentences into the dataset. During frequency analysis, we focus on sentences where the counterfactuals show no activation. In such cases, the target feature typically ranks among the top five, greatly enhancing search efficiency.
B.3 Multilingual Dataset
A base vector with robust representational capacity should activate in response to relevant sentences in multiple languages. We observed that many features activate only in one language, which is undesirable. Therefore, we constructed the “20-mix-2” dataset with a 1:1 ratio of English to Chinese sentences, ensuring that the extracted features possess cross-linguistic representational ability.
B.4 Weak Activations
In feature extraction, we observe that some features exhibit weak activation at “interference positions” where the target linguistic feature is not expected. These weak activations occur at positions that either match the cue words of the target feature, share similar morphology with the cue words, or belong to the same linguistic phenomenon category. The weak activations gradually diminish with increasing layer number until they vanish. This indicates that such weak activations are intermediate by-products of the model’s internal processing rather than being determined by the inherent connection strength between the base vector and the feature. Therefore, we exclude these weak activations (activations below one-fifth of the maximum activation value) when calculating the necessity probability of a base vector on the dataset. We present examples from morphology, syntax, and semantics as follows:
For the agentive suffix feature (12L248164), consider the following sentences:
(a) The journalist interviewed the mayor about the new policy. The “journalist” token has an activation value of 0.968.
(b) The newspaper published an article about the new policy. The “newspaper” token has an activation value of 0.197.
For the intransitive verb feature (17L63597), consider the sentences:
(a) She traveled to the supermarket. The “traveled” token has an activation value of 1.368.
(b) She drove the car to the supermarket. The “drove” token has an activation value of 0.362.
For the simile feature (26L75327), consider the sentences:
(a) Her eyes sparkled like stars in the night sky. The “like” token exhibits an activation value of 3.288.
(b) He looks like his father. The “like” token has an activation value of 0.557.
In these examples, the activations in the (b) sentences are weak and are not considered during feature extraction and evaluation. The presence of weak activation suggests that the model initially activates a broad range of potential semantics and then, during deeper processing, emphasizes the correct, contextually appropriate semantics. This observation warrants further investigation.
Appendix C Intervention Experiment Details
C.1 Intervention Cases
@@
Enhancement Can I textbfplease have your email address? The room should textbfbe large and well lit. It should textbfbe airy and bright and airy.
Default May I have your phone number? Sure, my ideal room has good ventilation and textbfis spacious.
@@
Enhancement "I was textbfasked for the story. " I having me textbfhad a “one the: ” textbftold. They: textbftold:
@@ C.2 LLM as a Judge
Appendix D Metric Calculation
D.1 Feature Representation Confidence (FRC)
@@
The first measure, the Probability of Necessity (PN), is defined as $PN=\frac{P(Y=1\mid do(X=1))-P(Y=1\mid do(X=0))}{P(Y=1\mid do(X=1))}$. This metric quantifies the extent to which the presence of a linguistic feature is necessary for the activation of a corresponding base vector. Here, $P(Y=1\mid do(X=1))$ represents the probability that the base vector is activated when the feature is present, whereas $P(Y=1\mid do(X=0))$ indicates the probability of activation when the feature is deliberately suppressed via intervention. The numerator, $P(Y=1\mid do(X=1))-P(Y=1\mid do(X=0))$, captures the net increase in activation due to the feature, and dividing by $P(Y=1\mid do(X=1))$ normalizes this increase relative to the activation when the feature is present.
Similarly, the second measure, the Probability of Sufficiency (PS) is expressed as $PS=\frac{P(Y=1\mid do(X=1))-P(Y=1\mid do(X=0))}{1-P(Y=1\mid do(X=0))}$. PS measures the likelihood that the introduction of the feature is sufficient to trigger the activation of the base vector. In this formulation, the denominator $1-P(Y=1\mid do(X=0))$ represents the maximum possible increase in activation probability (i.e., the probability that the base vector is not activated in the absence of the feature). Thus, PS reflects the proportion of this potential increase that is realized when the feature is present.
Finally, the Feature Representation Confidence (FRC) is computed as the harmonic mean of PN and PS: $FRC=\frac{2\,PN\,PS}{PN+PS}$. The harmonic mean is chosen because it ensures that FRC remains low if either PN or PS is low, thereby providing a balanced measure that only yields a high score when both necessity and sufficiency are strong. This approach allows us to robustly quantify the ability of the SAE latent space’s base vectors to represent the targeted linguistic features.
D.2 Feature Intervention Confidence (FIC)
@@
Appendix E Implementation Details

### v2

@@ Appendix A Dataset Construction
A.1 Dataset Description
The datasets are named according to the pattern “Feature Name+Feature Domain.” When a feature pertains to multiple linguistic domains, domains are concatenated with “&.” In total, the collection comprises 145 linguistic features, of which 99 are English features and 46 are Chinese features. Each feature-specific dataset contains 50 positive sentences and 50 counterfactual negative sentences.
A.2 Dataset Example
10-verbal_suffix-Morphology
He was able to stabilize the situation.
He was able to stable the situation.
The team has worked hard to solidify
their position in the market.
The team has worked hard to make their
position in the market solid.
43-copular_be-Syntax
My grandmother was a nurse.
My grandmother worked as a nurse.
Summer is the best season.
Summer ranks as the best season.
80-given_known-Pragmatics&Semantics
Have you seen the blue notebook anywhere?
Have you seen blue notebook anywhere?
That customer complained about service.
A customer complained about service.
111--&
130--&
A.3 Dataset Construction Guidelines
Work Content:
@@ A.3 Dataset Construction Guidelines
    For each linguistic feature, construct a dataset comprising 50 sentence pairs (100 sentences). Each pair contains one positive sentence and one negative sentence.
@@
    A positive sentence contains the target linguistic feature; a negative sentence is produced by minimally modifying its corresponding positive sentence so that it no longer contains that feature while preserving the smallest possible semantic difference and remaining grammatically correct (this operation is referred to as a “counterfactual” in causal analysis).
Notes:
  1. 1.
    Diversity: Ensure coverage of the feature’s common constructions and markers.
  2. 2.
    Counterfactual: Verify that the counterfactual edits are reasonable—including minimal change, human interpretability, and complete feature removal.
@@
    Ethical Check: Confirm that no sentence in the dataset contains discriminatory, biased, or harmful content.
@@
    Language-Specific Construction: Tailor construction to the particular characteristics of each language.
Specific Dataset Construction Process:
  1. 1.
    Manually create 5 sentences containing the feature, and for each, manually produce a counterfactual sentence—yielding 5 sentence pairs.
  2. 2.
    Expand these to 50 pairs using DeepSeek-R1 for Chinese and GPT-o4 for English, then apply manual edits guided by the Notes.
  3. 3.
    Conduct cross-review: volunteers who build the Chinese dataset review the English dataset, and vice versa, checking each item in the order specified under Notes.
Appendix B Feature Extraction Details
B.1 Feature Independence Validation
Condition Past-Tense Adversativity Intransitive Verb
Self 80/80 76/80 74/80
Control 1 -er 0/80 Sequential 0/80 Transitive Verb 0/80
Control 2 -ing 0/80 Causal 0/80 Ditransitive Verb 0/80
Control 3 -less 0/80 Parallel 0/80 Linking Verb 0/80
Control 4 -ness 0/80 Conditional 0/80 Modal Verb 0/80
Table 5: Activation ratios (activated/total) for target features and control conditions.
Sparse autoencoders (SAEs) effectively disambiguate neuron-level semantic polysemy, and this capability extends to representations of linguistic features.
We quantify feature independence using the necessity probability (PN) component of the Feature-Relevance Coefficient (FRC). PN measures the likelihood that a basis vector remains inactive when its associated feature is absent; a high PN therefore indicates that the vector is not spuriously activated by unrelated inputs, confirming its specificity to the intended phenomenon.
To further validate this independence, we evaluate each feature’s basis vector under multiple control conditions featuring superficially similar but semantically distinct constructions. Table 5 reports, for each feature, the ratio of sentences in which the vector activates (“activated/total”). Across all controls, activation rates are effectively zero, demonstrating that our selected basis vectors do not respond to non-target phenomena.
B.2 Feature Extraction Procedure
During feature extraction, we adhere to the following steps:
  1. 1.
    Input the feature-specific dataset into the model and encode each layer’s activations into a sparse latent space using Sparse Autoencoders (SAEs).
  2. 2.
    Compute the probability of sufficiency (PS), probability of necessity (PN), feature-relevance coefficient (FRC), and mean maximum activation for all basis vectors; then sort these vectors in descending order by FRC and select the top ten.
  3. 3.
    Employ a large-model agent to automatically analyze the activation patterns of the candidate basis vectors over the dataset, confirming their linguistic relevance to the target feature and characterizing their representational profiles.
  4. 4.
    For features undergoing further analytical or intervention experiments, manually review the basis vectors identified by the large-model agent to ensure the rigor of the experimental design.
B.3 Feature Extraction Prompt
We employ GPT-4o as the agent model for automated feature extraction. The system prompt is as follows:
Listing 1: Prompt for SAE Base-Vector Interpretation
You are an expert assistant for interpreting sparse
autoencoder (SAE) base vectors.
You will receive exactly one JSON object as input
with this structure:
"analysis_input": {
"layer": "00",
"base_vectors": [
"base_vector_id": 132317,
"tokens": ["The", "cat"],
"activations": [0.12, 0.05],
"ps": 0.62,
"pn": 0.58,
"frc": 0.60,
"avg_max_activation": 0.12
"base_vector_id": 81833,
"tokens": ["was", "chased"],
"activations": [0.08, 0.14],
"ps": 0.75,
"pn": 0.65,
"frc": 0.70,
"avg_max_activation": 0.14
"target_features": ["passive"]
Return exactly one JSON object with this schema:
"layer": "00",
"base_vectors": [
"base_vector_id": 132317,
"interpretation": "Marks passive voice
constructions",
"ps": 0.62,
"pn": 0.58,
"frc": 0.60,
"avg_max_activation": 0.12
"base_vector_id": 81833,
"interpretation": "Detects passive
participle forms",
"ps": 0.75,
"pn": 0.65,
"frc": 0.70,
"avg_max_activation": 0.14
"target_features": ["passive"]
Example 2:
Input:
"analysis_input": {
"layer": "08",
"base_vectors": [
"base_vector_id": 248593,
"tokens": ["runs"],
"activations": [0.45],
"ps": 0.76,
"pn": 0.96,
"frc": 0.85,
"avg_max_activation": 0.45
"base_vector_id": 62411,
"tokens": ["quickly"],
"activations": [0.32],
"ps": 0.82,
"pn": 0.90,
"frc": 0.88,
"avg_max_activation": 0.32
"target_features": ["adverbial_suffix"]
Output:
"layer": "08",
"base_vectors": [
"base_vector_id": 248593,
"interpretation": "Highlights adverbial
suffixes on verbs",
"ps": 0.76,
"pn": 0.96,
"frc": 0.85,
"avg_max_activation": 0.45
"base_vector_id": 62411,
"interpretation": "Detects adverbial
modifiers",
"ps": 0.82,
"pn": 0.90,
"frc": 0.88,
"avg_max_activation": 0.32
"target_features": ["adverbial_suffix"]
Requirements:
- Return only the JSON-no extra text.
- Round all floats to two decimal places.
- Preserve the input order of base_vectors.
- Echo layer and target_features exactly.
Appendix C Consistency Experiment
We introduce three linguistics experts to conduct an external consistency review of the GPT-4o proxy analysis. We sample 290 candidate base vectors and their activation patterns for the experiment. The consistency results are as follows:
Model Yes Model No Total
Human Yes TP = 120 FN = 1 121
Human No FP = 10 TN = 159 169
Total 130 160 290
Table 6: Consistency results comparing human judgments and model predictions.
Based on these results and further analysis of disagreement cases, human experts apply more flexible and lenient criteria under ambiguous activation patterns compared to GPT-4o, but overall consistency is very high—particularly the reliability of GPT-4o’s positive annotations.
Appendix D Intervention Experiment Details
D.1 Intervention Cases
@@
Enhancement Can I please have your email address? The room should be large and well lit. It should be airy and bright and airy.
Default May I have your phone number? Sure, my ideal room has good ventilation and is spacious.
@@
Enhancement "I was asked for the story. " I having me had a “one the: ” told. They: told:
@@ D.2 LLM as a Judge
Appendix E Metric Calculation
E.1 Feature Representation Confidence (FRC)
@@
The Feature Representation Confidence (FRC) is computed as the harmonic mean of PN and PS: $FRC=\frac{2\,PN\,PS}{PN+PS}$. The harmonic mean is chosen because it ensures that FRC remains low if either PN or PS is low, thereby providing a balanced measure that only yields a high score when both necessity and sufficiency are strong. This approach allows us to robustly quantify the ability of the SAE latent space’s base vectors to represent the targeted linguistic features.
E.2 Feature Intervention Confidence (FIC)
@@
Appendix F Linguistic Structure
F.1 Linguistics Levels
Morphology
The study of the internal structure of words—how roots, prefixes, suffixes, and inflectional endings combine to create different word forms and convey grammatical information such as tense, number, or case.
Syntax
The study of how words are arranged into larger units—phrases, clauses, and sentences—and the rules that govern their permissible order and hierarchical relationships within a language.
Semantics
The field that investigates meaning at the level of words, phrases, and sentences: how linguistic expressions map to concepts, objects, events, or states of affairs in the world, and how compositional principles let smaller meanings combine into larger ones.
Pragmatics
The study of how context and communicative intentions shape meaning in real-world use—how speakers choose utterances to achieve goals, how listeners infer implied or indirect meaning, and how factors like shared knowledge, discourse history, and social norms influence interpretation.
F.2 Linguistic Feature List
past_tense
Morphology & Semantics — verb form that locates an event before speech time.
noun_plural
Morphology — form marking more than one noun referent.
agentive_suffix
Morphology — suffix creating nouns for the doer of an action.
negation_prefix
Morphology — prefix that reverses or denies the base meaning.
degree_prefix
Morphology — prefix intensifying or scaling the base concept.
temporal_prefix
Morphology — prefix adding time relations such as “pre-” or “post-”.
quantitative_prefix
Morphology — prefix conveying amount or number.
spatial_or_directional_prefix
Morphology — prefix indicating place or direction.
nominal_suffix
Morphology — suffix that turns a base into a noun.
verbal_suffix
Morphology — suffix that turns a base into a verb.
adjectival_suffix
Morphology — suffix that turns a base into an adjective.
adverbial_suffix
Morphology — suffix that turns a base into an adverb.
possessive_form
Morphology & Syntax — morphological marking of ownership or relation.
third_person_singular
Morphology & Syntax — verb agreement form for he/she/it.
past_participle
Morphology & Syntax — verb form used in perfect aspect or passive voice.
present_participle
Morphology & Syntax — “-ing” form used for progressives or gerunds.
comparative
Morphology & Semantics — form showing a higher degree of a property.
superlative
Morphology & Semantics — form showing the highest degree of a property.
past_tense_irregular
Morphology — past form that does not end in “-ed”.
past_participle_irregular
Morphology — irregular past participle form.
intransitive_verb
Syntax — verb that takes no direct object.
transitive_verb
Syntax — verb that requires a direct object.
linking_verb
Syntax — verb that links subject to a complement.
anaphor
Syntax & Pragmatics — expression that refers back to an antecedent.
subject_auxiliary_inversion
Syntax — swapping subject and auxiliary (e.g., questions).
subject_verb_inversion
Syntax — reversing subject and main verb order.
passive_voice
Syntax & Semantics — clause where patient becomes grammatical subject.
subjunctive_mood
Syntax & Semantics — form expressing wish, doubt, or hypothetical state.
first_conditional
Syntax & Semantics — “if + present, will + verb” for real future possibility.
indirect_speech
Syntax & Pragmatics — reporting speech without a direct quote.
elliptical_sentences
Syntax — sentences with understood but omitted elements.
cleft_sentences
Syntax — “it + be + focus” construction for emphasis.
appositives
Syntax — noun phrase renaming another noun phrase.
non_defining_relative_clauses
Syntax — extra, non-restrictive relative clauses.
emphatic_structure
Syntax & Pragmatics — construction that highlights or stresses a clause part.
noun_clauses
Syntax — subordinate clauses functioning as nouns.
relative_clauses
Syntax — clauses that modify a noun with a relative word.
imperative_sentence
Syntax & Pragmatics — clause issuing a command or request.
of_genitive
Syntax — possession expressed with an “of” phrase.
s_genitive
Syntax — possession marked with apostrophe-s.
clausal_subjects
Syntax — clauses acting as the subject of a sentence.
extraposition
Syntax — moving a heavy subject/object to clause end with dummy “it”.
copular_be
Syntax — “be” used as a linking verb, not as an auxiliary.
echo_questions
Syntax & Pragmatics — repetition of prior utterance to seek confirmation.
tag_questions
Syntax & Pragmatics — short question tags appended to statements.
direct_object
Syntax — noun phrase receiving the verb’s action.
universal_quantifiers
Syntax & Semantics — words like “all, every” signifying totality.
existential_quantifiers
Syntax & Semantics — words like “some, any” signifying existence.
expletive
Syntax — syntactic placeholder such as “it” or “there”.
factives
Semantics & Syntax — predicates presupposing truth of their complement.
futurates
Semantics & Syntax — present-tense forms referring to scheduled future events.
intensifiers
Semantics & Pragmatics — adverbs that strengthen degree (e.g., “very”).
mass_noun
Syntax & Semantics — noun for uncountable substances (e.g., “water”).
object_expletives
Syntax — expletive pronouns occupying object position.
nominal_adverbials
Syntax — noun phrases functioning like adverbs.
split_infinitives
Syntax — placing a word between “to” and the verb stem.
quantifier
Syntax & Semantics — word or phrase expressing quantity.
count_nouns
Syntax & Semantics — nouns that can be enumerated individually.
active_verbs
Syntax — verbs used in active voice constructions.
middle_verb
Syntax & Semantics — verb whose subject is patient but appears active.
referring
Semantics & Pragmatics — linguistic act of pointing to real-world entities.
static_dynamic
Semantics — distinction between state verbs and action verbs.
punctual_durative
Semantics — contrast between instantaneous and durational events.
telic_atelic
Semantics — events with inherent endpoints vs. those without.
past
Semantics — temporal reference before the present moment.
future
Semantics — temporal reference after the present moment.
present_progressive
Semantics — aspect for ongoing present actions.
present_perfect
Semantics — aspect connecting past event to present state.
past_progressive
Semantics — aspect for ongoing past actions.
past_perfect
Semantics — event completed before a past reference point.
future_progressive
Semantics — ongoing action projected into the future.
future_perfect
Semantics — event completed before a future reference point.
epistemic
Semantics & Pragmatics — modality expressing speaker’s judgment of likelihood.
deontic
Semantics & Pragmatics — modality expressing obligation or permission.
spatial
Semantics — meaning elements relating to location or space.
person
Semantics & Pragmatics — grammatical category distinguishing speaker, addressee, others.
temporal
Semantics — meaning elements relating to time relations.
given_known
Pragmatics & Semantics — information already shared by speaker and listener.
representative
Pragmatics — speech act conveying assertions or descriptions.
directive
Pragmatics — speech act intended to get the hearer to act.
commisive
Pragmatics — speech act committing speaker to future action.
expressive
Pragmatics — speech act revealing speaker’s feelings or attitude.
declaration
Pragmatics — speech act that changes social reality.
metaphor
Semantics & Pragmatics — figurative transfer of meaning based on similarity.
synecdoche
Semantics & Pragmatics — figure where part stands for whole or vice versa.
non_synecdoche_metonymy
Semantics & Pragmatics — metonymic shift based on association, not part-whole.
coordination
Syntax & Semantics — joining of equal grammatical elements.
transitional
Semantics & Pragmatics — discourse element marking a shift or progression.
resultative
Syntax & Semantics — construction expressing a resultant state of an action.
optative
Syntax & Pragmatics — form expressing a wish or hope.
existential
Semantics & Syntax — clause asserting existence of something.
interrogative
Syntax & Pragmatics — clause type used for asking questions.
deixis
Pragmatics & Semantics — reference that depends on context (e.g., “here”, “now”).
turn_taking
Pragmatics — conversational management of who speaks when.
euphemism
Pragmatics & Semantics — mild term replacing a harsher one.
personification
Semantics & Pragmatics — giving human traits to non-human entities.
hyperbole
Semantics & Pragmatics — deliberate exaggeration for effect.
discourse_markers
Pragmatics — words that organize or signal discourse flow.
politeness
Pragmatics — linguistic strategies that mitigate imposition or face threat.
性_抽象名词后缀
形态学 — 后缀 “-性” 构成表示 “-ness/-ity” 的抽象名词。
化_动词性后缀
形态学 — 后缀 “-化” 构成动词,表示“使…/成为…”。
们_复数后缀
形态学 & 语义学 — 后缀 “-们” 标记人称复数。
重叠构词
形态学 & 语义学 — 通过词素重叠构词,以强调或表迭代。
不及物动词
句法学 & 语义学 — 不能带直接宾语的动词。
及物动词
句法学 & 语义学 — 需要直接宾语的动词。
系动词
句法学 — 连接主语与补语的动词。
属格
句法学 & 语义学 — 所有格或所属关系的语法标记。
逆向结构
句法学 & 语义学 — 为强调或疑问而颠倒正常语序。
被动语态
句法学 & 语义学 — 将承事者作为句法主语的被动结构。
主题_述评句
句法学 & 语用学 — 将句子拆分为主题和述评部分的结构。
回指
句法学 & 语义学 & 语用学 — 指代先行项的表达方式。
间接引语
句法学 & 语用学 — 不引用原话的转述形式。
省略句
句法学 & 语用学 — 上下文可恢复的省略结构。
同位结构
句法学 — 两个等价名词短语并列重命名的结构。
反问句
句法学 & 语用学 — 期望无真实答案的修辞性疑问句。
感叹词
语用学 — 表达突发情感的独立词。
祈使句
句法学 & 语用学 — 用于发布命令或请求的句式。
语气助词
形态学 & 语义学 & 语用学 — 表示说话人态度的助词。
轻动词
句法学 & 语义学 — 与名词搭配使用,语义轻的动词。
主观数量
语义学 & 语用学 — 说话人评估的模糊数量表达。
使役结构
句法学 & 语义学 — 表示“使/让某人做…”的致使结构。
条件句
句法学 & 语义学 — 表达“如果…,就…”条件关系的句子。
兼语句
句法学 — 一个名词在结构中既作宾语又作主语。
情态
语义学 & 语用学 — 表示能力、必要性等的情态范畴。
时体标记
形态学 & 语义学 — 标记时态或体的形式。
假设
语义学 & 语用学 — 表示假设情景的表达。
受事主语句
句法学 & 语义学 — 主语为动作承事者的句子。
可能
语义学 & 语用学 — 表示可能性或潜在性的表达。
因果
语义学 & 语用学 — 表示因果关系的表达。
并列
句法学 & 语义学 — 平等地并列元素的结构。
明喻
语义学 & 语用学 — 用“像”等词显性标记的比喻。
暗喻
语义学 & 语用学 — 无显性比较词的隐喻。
比较
语义学 — 表示相似或差异的语言表达。
致使
句法学 & 语义学 — 表示结果状态的致使表达。
让步
语义学 & 语用学 — 虽承认…但仍…的让步关系。
转折
语义学 & 语用学 — 标记对比或转折的关系。
递进
语义学 & 语用学 — 表示进一步增强信息的关系。
指示
语义学 & 语用学 — 根据上下文指示实体的表达。
话轮转换
语用学 — 对话中管理轮到谁发言的结构。
委婉语
语用学 — 缓和直接性的委婉表达。
拟人
语义学 & 语用学 — 将人类特征赋予非人实体的表达。
夸张
语义学 & 语用学 — 为强调而故意夸大的表达。
话语标记
语用学 — 引导和组织话语流程的词语。
礼貌
语用学 — 表示礼貌或维护面子策略的语言手段。
数量词
句法学 & 语义学 — 数词加量词短语,表示确切数量。
Appendix G Implementation Details